OpenStudy (anonymous):

Calculate the average atomic mass of rubidium. Rubidium has two isotopes, 85^Rb and 87^Rb. 85^Rb has an atomic mass of 84.912 amu and occurs at an abundance of 72.17%. 87^Rb has an atomic mass of 86.909 amu and occurs at an abundance of 27.83%. Show your work.

OpenStudy (frostbite):

To try calculate the atomic masses of our molecules we define calculate the relative (average) atomic masses. The relative/average comes from the idea that we weight the masses with their probability of occurrence. In terms of math: \[\large M_{element}=\sum_i m_i \times p_i\] Here is \(m_i\) the mass of the \(i\)'th isotope with the probability of occurrence \(p_i\). If the math confuses you, I can show you more specifically how to apply this for rubidium.

OpenStudy (anonymous):

Yeah if you don't mind showing the math to me that would be really helpful

OpenStudy (frostbite):

Basically the function just say we must add up these "weighted masses" So: \[\large M_{ribidium}=m(^{85}\text{Rb}) \times p(^{85}\text{Rb}) +m(^{87}\text{Rb}) \times p(^{87}\text{Rb})\] From here it is all about plugging in numbers, just remember to make your percentages into decimal numbers.

OpenStudy (anonymous):

Thank you so much for your help! :)

OpenStudy (frostbite):

No problem at all. If you want to write out your answer I can check it for you before you turn it in.